OpenStudy (anonymous):

haemoglobin contains 0.33% of Fe by wt. The molecular mass of haemoglobin is approximately 67200. The no. of Fe atoms ( at wt. of Fe 56) present in one molecule of Haemoglobin is a) 6 b) 1 c) 4 d) 2 P.S. I don't want to get the answer or the formula ( i know them) I need to know the procedure of doing such questions :)

OpenStudy (aaronq):

so what you have (0.33%) is the mass percent composition of iron. What you wanna do if find the total weight (mass) of iron in the sample. Then use good old \(n=\dfrac{m}{M}\) to find moles. then use: \(n=\dfrac{N}{N_A}\) to find the number of atoms present.

OpenStudy (anonymous):

do you mean to say that i need to find out \[Fe = .33% *67200 \] gms and then divide that by 56 ( atomic wt of Fe) ?

OpenStudy (anonymous):

Fe = .33% * 67200

OpenStudy (aaronq):

yeah you need to find the mass of iron in the hemoglobin molecule. yep thats it.

OpenStudy (aaronq):

wait i misread the question. It's asking for the number of atoms in 1 Hb. so you should divide the molecular mass of Hb by avogadros number first.

OpenStudy (anonymous):

that means 221.76/56 = 4 moles :) Thanks buddy!

OpenStudy (aaronq):

which means 4 atoms per Hb

OpenStudy (anonymous):

so that means it 4 Atoms per Hb.! not 4 moles!!!!

OpenStudy (aaronq):

yes, thats why i said you had to divide the MW by avogadros number (to find the mass of 1 Hb molecule, not a whole mole).

OpenStudy (anonymous):

Alright! Thanks